# Technical Note: Vendor-Specific Approach for Standardized Uptake Value Calculation

### v1

@@ Abstract
Purpose: This technical note introduces a robust, vendor-specific SUV calculation strategy that addresses inconsistencies in current methodologies. We also integrate this strategy into an open-source software solution, Z-Rad, capable of converting raw PET DICOM data into body-weight normalized SUV NIfTI files.
Methods: Our SUV calculation strategy was developed by reviewing DICOM conformance statements from GE, Philips, and Siemens. Validation was conducted using real-world PET datasets, and the proposed strategy was compared to existing software solutions.
Results: Our SUV calculation approach demonstrated improved accuracy, particularly in resolving time-related discrepancies in the studied data. Our analysis also identified inconsistencies in the SUV calculation methods used by popular commercial and open-source software solutions, which do not fully account for vendor-specific DICOM attributes and PET image acquisition times. These limitations resulted in errors in SUV estimation reaching 33% when comparing our strategy to studied software.
Conclusions: The proposed vendor-specific SUV calculation strategy significantly enhances accuracy in PET imaging by addressing key inconsistencies caused by variations in DICOM attributes and image acquisition times across different vendors. This method effectively reduces SUV calculation errors and has been integrated into an open-source software, Z-Rad.
@@
1 Introduction
The standardized uptake value (SUV) is a fundamental metric in positron emission tomography (PET) 1, shown to be predictive for several tumor types 2, 3, essential for quantifying treatment response4, and is a precursor to more comprehensive metrics like total lesion glycolysis (TLG) or metabolic tumor volume (MTV) 5. However, the methods for calculating SUV from DICOM files vary across vendors 6, 7, 8, software solution9, and open-source datasets 10.
A significant contribution to the standardization of SUV calculations was made by Paul Kinahan et al. 11, who surveyed MIM, Philips, Siemens, and GE, and derived both vendor-specific 12 and vendor-neutral 13 SUV calculation strategies. Their vendor-specific approach 12 takes into account private DICOM attributes from different vendors, but does not completely account for the errors that occur when the Series Time (0008, 0031) is altered. In contrast, their vendor-neutral "happy path only" strategy 13 ignores all vendor-specific DICOM attributes and relies solely on Series Time (0008, 0031).
In this technical note, we present an adapted and comprehensive vendor-specific strategy for SUV calculation, addressing various pitfalls and challenges associated with real-world PET data. We provide a thorough analysis of existing SUV calculation methodologies, highlight potential issues, and propose a software solution to improve comparability of inter-vendor PET data.
@@ 2.1 SUV Definition
The general equation for the SUV normalized by the body weight ($SUV_{bw}$) is defined as follows:
$\text{SUV}_{bw}=\frac{A\cdot w}{D_{i}}\cdot 2^{\frac{t_{c}-t_{i}}{T_{1/2}}}\quad\left[\frac{g}{ml}\right],$ (1)
where $D_{i}$ is the injected dose of the radiotracer (in Bq), $w$ patient’s weight (in g), $A$ activity concentration (in Bq/mL), $t_{c}-t_{i}$ difference between measurement and radiotracer injection times respectively (in s), and $T_{1/2}$ radiotracer half-life time (in s).
@@ 2.2 Calculation of SUV from DICOM
In PET DICOM files, the Rescale Slope (0028,1053) and Rescale Intercept (0028,1052) are used to convert the pixel values stored in the image into physical units, as defined by Units (0054,1001). These units can represent measurements such as activity concentration in Becquerels per milliliter (BQML), grams per milliliter (GML), or counts (CNTS). This conversion is crucial because the raw pixel values directly from the scanner do not represent the radiotracer concentration. A useful parameter for further discussion, $A^{\prime}$, can be derived from the DICOM file using the following equation:
@@
where $P$ represents the original stored pixel value, $S$ is the Rescale Slope (0028,1053), and $I$ is the Rescale Intercept (0028,1052). The parameter $A^{\prime}$ is measured in the units specified in Units attribute. If units are equal to BQML, then $A^{\prime}$ is equal to $A$ from Eq. 1.
@@
Most discrepancies in SUV calculation from DICOM files originating from different vendors arise from different definitions of the measurement time $t_{c}$ (Eq. 1) and incorrect handling of different PET DICOM units. To address these issues, we derived a vendor-specific SUV calculation strategy based on the DICOM conformance statements from three major PET machine manufacturers: Philips 7, Siemens 8, and GE 6. Leveraging this information, we developed a unified framework to ensure consistency and accuracy in SUV definition across different systems (Fig. 1). A comprehensive list of all used DICOM attributes is provided in the Appendix.
2.2.1 Decay Correction Time
@@
In metabolic PET imaging, the most common Units (0054, 1001) value is BQML. For this unit, Siemens, Philips, and GE agree that when the Decay Correction (0054, 1102) is set to ADMIN all images are decay-corrected to the tracer injection time $t_{i}$. This implies that $t_{c}$ is equal to $t_{i}$ in Eq. 1.
If the Decay Correction (0054, 1102) is START, vendors provide private DICOM attributes indicating the specific decay correction time $t_{c}$ to which the image is corrected. For instance, Siemens uses the private tag (0071, 1022), and GE uses (0009, 100d). If the corresponding tag is not present, or if the vendor is Philips - which does not have a dedicated private attribute - and the DICOM series includes multiple acquisition times, the decay correction time $t_{c}$ is defined as the earliest acquisition time in the series.
@@ 2.2.2 PET DICOM Units
Philips machines can also generate PET data with units of CNTS (counts). To calculate the SUV or the activity concentration (in Bq/mL) in this case, we refer to the equations provided in the Philips DICOM conformance statements 7:
@@
Another commonly used PET DICOM unit is GML (grams per milliliter). Although Siemens, GE, and Philips do not mention this unit in their DICOM conformance statements, the DICOM standard documentation 14 indicates that if the unit is GML and the SUV Type (0054, 1006) is set to BW (body-weight) or left unspecified, the SUV should be interpreted as a body-weight normalized and equal to $A^{\prime}$ (Eq. 2).
@@ 2.3 Validation datasets and Evaluated Software Solutions
To evaluate the proposed SUV calculation strategy (Fig. 1) and identify potential pitfalls, we tested it using PET datasets from our institution and from the autoPET challenge 10, which include metabolic PET data from Siemens, GE, and Philips (Table 1). Our strategy has been implemented in the open-source software Z-Rad 15, which can be used to convert raw PET DICOM files into body-weight normalized SUV NIfTI files. Additionally, we reviewed third-party software tools, such as the open-source software 3D Slicer and the commercial software MIM. Both solutions are widely used for reading PET DICOM files. This assessment helped us highlight any inconsistencies that may occur in their SUV calculations.
@@ 3 Results
A comparison of SUV values between Z-Rad, MIM, and 3D Slicer for GE data showed agreement among all three software for 59% of the images. For the remaining GE data, there was an average error of 31% for 3D Slicer and 32% for MIM resulting in underestimation of the SUV. This was caused by the image series time being shifted one hour earlier than the earliest acquisition time. In these cases, the GE private tag (0009, 100d) contained the correct decay correction time, which matched the earliest series acquisition time. MIM was able to detect when the series time is earlier than the injection time (16.5% of the data) and such images were not converted to SUV. However, MIM did not report errors when the series time was later than the injection time but still earlier than the earliest acquisition time (24.5% of the data). 3D Slicer filed to detect either of these two data related issues.
For Siemens data, the software solutions agreed for almost all images, except for three cases where a time difference of over an hour between the series and the earliest acquisition times resulted in an average SUV underestimation error of 33%. Another issue identified in the Siemens data was that the private Decay Correction Date Time attribute (0071, 1022) had the correct time but an incorrect date. Despite the series and acquisition dates being consistent, this could potentially lead to an erroneous addition of 86’400 seconds (24 hours) to $t_{c}$ (Eq. 1).
All Philips scans with BQML units showed a slight inconsistency between the series time and the earliest acquisition time. However, the maximum error in SUV estimation did not exceed 1.5%. For Philips images with CNTS, 3D Slicer’s PET module 9 did not apply the Philips-specific conversion factors (Eq. 3, Eq. 4), instead attempting to convert the data based on the Series Date and Series Time attributes. Similarly, MIM did not convert Philips PET DICOM files from CNTS to BQML, displaying the raw image in CNTS, which prevented any SUV-based clinical conclusions from being drawn. For the studied PET scans with CNTS, Philips provided private DICOM attributes that allow to convert image intensity to SUV (Eq. 3). However, since the SUV Type (0054,1006) was not specified in studied scans, we verified that Eq. 3 lead to SUV normalized to patient’s body weight in studied data.
@@ 4 Discussion
The vendor-neutral “happy path only” strategy proposed by Paul Kinahan et al. 13 assumes that $t_{c}$ is equal to the DICOM series time. However, this assumption is not always true. As shown in our results, this approach leads to errors when there is a mismatch between the earliest acquisition time and the series time, or when vendor-specific DICOM attributes are needed. The “happy path only” strategy is used in 3D Slicer, as stated in its documentation 9. While we do not know the specifics of SUV estimation methodology in MIM, our results suggest that it follows a similar “happy path only” vendor-neutral strategy, as it neither incorporated any vendor-specific DICOM attributes from the studied data nor correctly estimated $t_{c}$.
Another example of incorrect SUV estimation is the strategy used in the autoPET challenge 10. This approach relied on the acquisition time attribute of individual slices, resulting in an error that increased over the imaging duration. For DICOM data where the Series Time attribute remained unaltered, the error began at 0% when the acquisition time matched the series time but could grow to 21% for the last slices due to a 30-minute difference between the earliest and latest acquisition times. This led to images that failed to accurately depict SUV differences between body parts, as the error was not constant but gradually increased with each consecutive slice.
A closer examination of the vendor-specific strategy proposed by Paul Kinahan et al. 12 revealed that it addresses potential pitfalls in Philips CNTS data and accounts for GE and Siemens private tags, but only when the series time is later than the acquisition time. However, this approach does not apply to cases where the series time is earlier. As a result, this discrepancy leads to significant errors similar to those seen with the vendor-neutral "happy path only" strategy, resulting in average errors of up to 33% in the studied data.

### v2

@@ Abstract
Purpose: This technical note introduces a robust, vendor-specific SUV calculation strategy designed to addresses inconsistencies in current methodologies. We also integrate this strategy into an open-source software, Z-Rad, capable of converting raw PET DICOM data into body-weight normalized SUV NIfTI files.
Methods: Our SUV calculation strategy was developed by reviewing DICOM conformance statements. Validation was conducted using real-world PET datasets, and the proposed strategy was compared to existing commercial and open-source software tools.
Results: Our proposed strategy demonstrated improved accuracy by resolving time-related discrepancies in the studied data, which can lead to substantial errors in SUV calculation. In particular, we identified that popular software solutions often overlook vendor-specific DICOM attributes and incorrect PET image series times, resulting in SUV estimation errors of up to 32% in our test data.
Conclusions: The proposed vendor-specific SUV calculation strategy facilitates quantitative PET imaging analysis by addressing variations in DICOM attributes encoding across different vendors. This method allows to avoid common mistakes with adherence to the DICOM conformance statements and has been integrated into an open-source software, Z-Rad.
@@
The standardized uptake value (SUV) is a fundamental quantitative metric in positron emission tomography (PET) Thie1431 to measure the relative concenration of a radiotracer in a specific region of interest providing a standardized way to assess metabolic activity. It has been shown to be predictive for several tumor types Lai2024, Shalaby2022, essential for quantifying treatment responseVanderhoek1188, and is a precursor to more comprehensive metrics like total lesion glycolysis (TLG) Kitao2016, metabolic tumor volume (MTV) Kitao2016, and PET-based radiomic studies. However, the methods for calculating SUV from DICOM files vary across vendors ge_healthcare_pet_dicom, philips_dicom_nuclear_medicine, siemens_dicom_molecular_imaging, software solutionslicer_suv_calculator, and online challenges autopet_dataset.
A major step toward standardizing SUV calculations was provided by Quantitative Imaging Biomarkers Alliance (QIBA) qiba_suv. Paul Kinahan et al. surveyed four different vendors to develop both vendor-specific qiba_suv_pseudocode and vendor-neutral qiba_suv_happypath QIBA SUV calculation strategies. Their vendor-specific approach qiba_suv_pseudocode takes into account private DICOM attributes from different vendors but only addresses errors when the Series Time (0008,0031) is after the Acquisition Time (0008,0022). In contrast, their vendor-neutral “happy path only” strategy qiba_suv_happypath forgoes all vendor-specific tags and relies solely on Series Time (0008,0031).
In this technical note, we present a refined, comprehensive vendor-specific SUV calculation strategy that addresses various pitfalls and challenges associated with real-world PET data. We provide a thorough analysis of existing SUV calculation methodologies, highlight potential issues, and propose a software solution to facilitate consistent, inter-vendor comparisons in PET imaging. Our objective is to improve quantitative reliability across a diverse range of DICOM datasets and scanning protocols.
@@ 1.1 SUV Definition
The standardized uptake value (SUV) normalized by the body weight, $\text{SUV}_{bw}$, is typically defined as:
$\text{SUV}_{bw}=\frac{A\cdot w}{D_{i}}\cdot 2^{\frac{t_{s}-t_{i}}{T_{1/2}}}\quad\left[\frac{g}{mL}\right],$ (1)
where $D_{i}$ is the injected dose of the radiotracer (in Bq), $w$ - patient’s weight (in grams), $A$ - activity concentration (in Bq/mL), $t_{s}-t_{i}$ - nonnegative difference between measurement and radiotracer injection times respectively (in seconds), and $T_{1/2}$ - radiotracer half-life time (in seconds).
@@ 1.2 Calculation of SUV from DICOM
1.2.1 Conversion of Stored Pixel Values to Physical Units
In PET DICOM files, the Rescale Slope (0028,1053) and Intercept (0028,1052) are used to convert the pixel values stored in the image into physical units, as defined by Units (0054,1001). These units can represent measurements such as activity concentration in Becquerels per milliliter (BQML), grams per milliliter (GML), or counts (CNTS). A useful parameter for further discussion, $A^{\prime}$, can be derived from the DICOM file using the following equation:
@@ 1.2.1 Conversion of Stored Pixel Values to Physical Units
where $P$ represents the original stored pixel value, $S$ is the Rescale Slope (0028,1053), and $I$ is the Rescale Intercept (0028,1052). The parameter $A^{\prime}$ is measured in the units specified in Units (0054,1001). If units are BQML, then $A^{\prime}$ is equal to $A$ from Eq. 1.
@@
Most discrepancies in SUV calculation from DICOM files arise from incorrect handling of PET DICOM units by third-party software and differences in decay correction strategies. To address these issues, we derived a vendor-specific SUV calculation workflow based on the DICOM conformance statements from three major PET machine manufacturers: Philips philips_dicom_nuclear_medicine, Siemens Healthineers siemens_dicom_molecular_imaging, and GE ge_healthcare_pet_dicom. Leveraging this information, we developed a unified framework to ensure consistency and accuracy in SUV definition across different systems (Figure 1). A comprehensive list of all used DICOM attributes is provided in the Supplement.
1.2.2 Scan Start Time Estimation
@@
In metabolic PET imaging, the most commonly used Units (0054,1001) value is BQML. For this unit, when the Decay Correction (0054,1102) is set to ADMIN, all images are decay-corrected to the radiopharmaceutical injection time $t_{i}$. This implies that $t_{s}$ is equal to $t_{i}$ in Eq. 1.
If the Decay Correction (0054,1102) value is set to START, the image frames are decay-corrected to the scan start time. In principle, this should align with Series Time (0008,0031). However, because Series Time can be altered by subsequent processing, Siemens Healthineers and GE provide private DICOM tags ((0071,1022) for Siemens Healthineers and (0009,100d) for GE) to store a more reliable Decay Correction Date Time.
If these private tags are missing, according to the GE DICOM Conformance Statement, the scan start time $t_{s}$ can be reconstructed as followingge_healthcare_pet_dicom:
$t_{s}=t_{a}-\Delta t_{FRT},$ (3)
where $t_{a}$ is the Acquisition Time (0008,0022) and $\Delta t_{FRT}$ is the Frame Reference Time(0054,1300) value in seconds. This indicates, that the images are decay corrected to the start of a frame. Similarly for Siemens Healthineers, the scan start time can be calculated based on the information from their DICOM Conformance Statement. Specifically, the provided definition of the Decay Factor implies that images are decay corrected to the time at which the average activity in a frame is corrected ($\Delta t_{AVE}$)siemens_dicom_molecular_imaging:
$t_{s}=\Delta t_{a}+\Delta t_{AVE}-\frac{\ln(\kappa)}{\lambda},$ (4)
$\Delta t_{AVE}=\frac{1}{\lambda}\ln{\frac{\lambda\Delta t_{AFD}}{1-e^{-\lambda\Delta t_{AFD}}}},$ (5)
where $\Delta t_{AFD}$ is the Actual Frame Duration (0018,1242) value (in seconds), $\kappa$ is the Decay Factor (0054,1321), and $\lambda$ = $\ln(2)/T_{1/2}$ is the decay constant, with $T_{1/2}$ specified in the Radiopharmaceutical Information Sequence (0054,0016). Since Siemens Healthineers implementation also sets Frame Reference Time(0054,1300) from the scan start to $\Delta t_{AVE}$, the equation:
$t_{s}=t_{a}+\Delta t_{AVE}-\Delta t_{FRT}$ (6)
can be used (confirmed by Siemens Healthineers). The same approach has been also suggested by Paul Kinahan et al. in the vendor-specific workflowqiba_suv.
Philips does not provide a dedicated private attribute for Decay Correction Date Time. Instead, for datasets where Decay Correction (0054,1102) equals START, Philips sets Decay Factor (0054,1321) to a placeholder value of 1.0 philips_dicom_nuclear_medicine. Thus, the effective scan start time must be derived using the Frame Reference Time (0054,1300) along with the “average time” $\Delta t_{AVE}$ for a decaying radionuclide (Eq. 6).
Bringing together equations 3, 5, 4, and 6 allows to derive a general equation for estimation of the scan time:
$t_{s}=\begin{cases}t_{a}+\frac{1}{\lambda}\ln\left(\frac{\lambda\Delta t_{AFD}}{1-e^{-\lambda\Delta t_{AFD}}}\right)-\frac{\ln(\kappa)}{\lambda}&\text{for Siemens Healthineers},\\[10.0pt] t_{a}-\Delta t_{FRT}&\text{for GE},\\[10.0pt] t_{a}+\frac{1}{\lambda}\ln\left(\frac{\lambda\Delta t_{AFD}}{1-e^{-\lambda\Delta t_{AFD}}}\right)-\Delta t_{FRT}&\text{for Philips and}\\ &\text{Siemens Healthineers.}\end{cases}$ (7)
@@ 1.2.3 PET DICOM Units
Although BQML is the most common in PET imaging, alternative units also appear in practice. Philips machines can also generate PET data with units of CNTS (counts). To calculate the SUV or the activity concentration (in Bq/mL) in this case, we refer to the equations provided in the Philips DICOM conformance statements philips_dicom_nuclear_medicine:
@@
Another commonly used PET DICOM unit is GML (grams per milliliter). Although Siemens Healthineers, GE, and Philips do not mention this unit in their DICOM conformance statements, the DICOM standard documentation dicom_part03_section_c8_9 indicates that if the unit is GML and the SUV Type (0054,1006) is set to BW (body-weight) or left unspecified, the SUV should be interpreted as a body-weight normalized and equal to $A^{\prime}$ (Eq. 2).
@@ 1.3 Validation datasets and Evaluated Software Solutions
To evaluate the proposed SUV calculation strategy (Fig. 1) and identify potential pitfalls, we tested it using PET datasets avaliable at our institution, as well as data from autoPET challenge autopet_dataset, gatidis_kuestner_2022, IBSI II validation data vallieres_2015_tcia, vallieres_2015_pmb, clark_2013_tcia, and ACRIN 6685 dataset Kinahan2019 which include metabolic PET data from Siemens Healthineers, GE, and Philips (Table 1). Our strategy has been implemented in the open-source software Z-Rad zrad, which can be used to convert raw PET DICOM files into body-weight normalized SUV NIfTI files. Also, we published our vendor-specific method for SUV calculation on GitHub Gist GitHubGist. Additionally, we reviewed third-party software tools, such as PET extentionPETDICOM_Slicer (version 6840644) for the open-source software 3D Slicer and the commercial software MIM (version 7.4.2). Both solutions are widely used for reading PET DICOM files. This assessment helped us highlight any inconsistencies that may occur in their SUV calculations.
@@ 2 Results
2.1 Validation of the Decay Correction Time Estimation
Vendor-specific private tags were present in 1,301 out of 1,658 BQML PET series from Siemens Healthineers and GE, corresponding to 485,864 out of 578,357 DICOM slices.
For GE data, private tag (0009,100d) was present in 130’717 DICOM slices. Eq. 3 was validated for all these slices, with no discrepancies observed.
For Siemens Healthineers data, the private tag (0071,1022) was present in 355’147 DICOM slices, allowing validation of both Siemens Healthineers-specific equations (Eq. 4 and Eq. 6). Eq. 4 was accurate within an absolute discrepancy of less than 1.0 s for 95% of slices, less than 2.0 s for 99.98% of slices, and did not exceed 3.8 s for the remaining slices. Validation using Eq. 6 (based on information from Siemens Healthineers) showed an absolute discrepancy below 1.0 s for 99.994% of slices.
Philips does not provide a private tag for decay correction date and time. Therefore, validation was performed by comparing results from Eq. 6 against Series Time (0008,0031). For the studied Philips data, the Series Time closely matched the earlierst Acquisition Time value, suggesting that the series time had not been modified. Validation on 9’136 DICOM slices revealed that the maximum deviation from Series Time was 1.5 s. For 99.1% of slices, the absolute discrepancy was below 1.0 s. Deviations exceeding 1.0 s occurred only in slices belonging to a single PET series.
2.2 Comparison of Z-Rad, MIM, and 3D Slicer
A comparative analysis of SUV values among Z-Rad, MIM, and 3D Slicer for the data in BQML units demonstrated that in 86% of scans, all three software tools yielded the same SUV values. However, for the remaining 14% (215 scans), lower SUV values were observed in 3D Slicer and MIM, with an average difference of 32%. Further investigation revealed that this difference came from the use of the general Series Time (0008,0031) rather than the vendor-specific private tags by 3D Slicer and MIM. In these 215 scans, the Series Time (0008,0031) was shifted by one hour relative to the information stored in the vendor-specific private tag.
MIM and 3D Slicer detected neither when the Series Time (0008,0031) was earlier than the injection time (in 86 cases) nor when it was later than the injection time but still earlier than time in vendor-specific private tag (or than the earliest Acquisition Time (0008,0032), when no private tag was present), which occurred in 126 cases, as a result producing lower SUV values. This was mostly eliminated by using the vendor-specific private tags (0009,100d) or (0071,1022). In cases where this private tag was not present, $t_{s}$ could be reliably calculated using Eq. 7. However, in 65.6% of the BQML data, private tag contained an shifted date compared to the Series Date (0008,0021) and Acquisition Date (0008,0022), which were consistent.
For the GML data, MIM handled SUV calculation by following the strategy outlined in the DICOM standard documentation dicom_part03_section_c8_9. In contrast, 3D Slicer applied the QIBA “happy path only” strategy qiba_suv_happypath, resulting in different SUV conversion.
For CNTS data, the 3D Slicer PET module slicer_suv_calculator did not apply the required vendor-specific conversion factors (Eqs. 8 and 9), relying instead on Series Time (0008,0031), while MIM did not convert PET DICOM files from CNTS to BQML and displayed images only in raw CNTS units.
The studied PET CNTS scans, provided private DICOM attributes enabling accurate SUV calculation via (Eq. 8). However, since the SUV Type attribute was not specified, we validated that Eq. 8 indeed produces SUV values normalized to patient body weight. This was verified for the data where both SUV Scale Factor (7053,1000) and Activity Concentration Scale Factor (7053,1009) were present.
2.3 Comparison of Proposed and QIBA SUV Calculation Strategies
When compared to our strategy, the QIBA “happy path only” approach qiba_suv_happypath accounts nither for alterations in time-related DICOM metadata nor for vendor-specific conversion rules. By design, the “happy path only” strategy applies to PET data in BQML units only, providing no support for other unit types. In contrast, our strategy supports vendor-specific conversion rules and BQML, GML, and CNTS units. Due to no additional handling of time-related DICOM metadata by the “happy path only” approach, it led to discrepancies in SUV values of 32% in 215 PET scans with BQML units from our validation dataset (Table 1).
The vendor-specific QIBA strategy proposed by Paul Kinahan et al. qiba_suv_pseudocode extends the “happy path only” approach by incorporating some vendor-dependent considerations but fails for GE data when early bed positions are missing. This limitation arises because it uses Eq. 6, whereas GE data require Eq. 3. In addition, this method only addresses cases where Series Time (0008,0031) is later than Acquisition Time (0008,0032), neglecting instances where the Series Time is earlier. This results in discrepancies identical to those of the QIBA “happy path only” approach — 32% in 215 PET scans—whereas our strategy is directly based on vendor-specific metadata and remains robust to general DICOM attributes alternations.
@@ 3 Discussion
Both QIBA strategies proposed by Paul Kinahan et al. qiba_suv_happypath, qiba_suv_pseudocode do not fully account for vendor-specific details and alterations in time-related DICOM metadata, leading to discrepancies in SUV values of up to 32% in the studied dataset. Therefore, it is recommended to use the time component indicated in private tags whenever possible, and when these tags are absent, to calculate the scan start time based on Acquisition Time (0008,0032), Decay Factor (0054,1321), Actual Frame Duration (0018,1242), and Frame Reference Time (0054,1300) as shown in Eq. 7. The proposed strategy combines vendor-specific attributes with fallback calculations based on general DICOM attributes when necessary, this method provides a robust and transparent solution to a persistent challenge in quantitative PET imaging. During the vendor-specific validation of the decay correction time estimation, we did not observe differences between models originating from the same vendor, indicating that all included vendors were consistent in their strategies across different PET scanner models.
Unfortunately, we were not able to obtain PET data from other vendors (e.g., Canon, United Imaging), thus limiting validation of the current strategy to scanners from Philips, GE, and Siemens Healthineers. Future work will include maintaining the proposed strategy and updating it according to vendor-specific DICOM conformance statements, as well as incorporating additional PET hardware vendors.
The proposed strategy is implemented in publicily available and open-source software Z-Rad zrad. Additionally, we provide lightweight Python code for SUV calculation on GitHub Gist GitHubGist.